Determine the displacement of a man if he walks 3 miles north then turns right and walks 4
miles east.
1
Expert's answer
2019-09-12T10:37:43-0400

Solution. Sketch a figure of the movement of a man




Find the magnitude of the displacement of men using the Pythagorean theorem.


"d^2=3^2+4^2=9+16=25"

"d=\sqrt 25=5"

Answer. 5 miles.